# A practical identifiability criterion leveraging weak-form parameter estimation

### v2

@@ Abstract
In this work, we define a practical identifiability criterion, $(e,q)$-identifiability, based on a parameter $e$, reflecting the noise in observed variables, and a parameter $q$, reflecting the mean-square error of the parameter estimator. This criterion is better able to encompass changes in the quality of the parameter estimate due to increased noise in the data (compared to existing criteria based solely on average relative errors). Furthermore, we leverage a weak-form equation error-based method of parameter estimation for systems with unobserved variables to assess practical identifiability far more quickly in comparison to output error-based parameter estimation. We do so by generating weak-form input-output equations using differential algebra techniques, as previously proposed by Boulier et al [1], and then applying Weak form Estimation of Nonlinear Dynamics (WENDy) to obtain parameter estimates. This method is computationally efficient and robust to noise, as demonstrated through two classical biological modelling examples.
@@ 1 Introduction
Parameter estimation is central to mathematical modeling in the biological sciences. However, despite the increasing sophistication of computational techniques over the years, there are many examples of researchers using the same model and obtaining widely varying parameter estimates [2, 3, 4, 5, 6]. Therefore, analyzing model identifiability, i.e., the ability to recover unique parameters of a system from observations, is critical to developing robust and replicable models of biological systems.
@@
Structural identifiability addresses whether model parameters can be uniquely determined given the observation of some output variables of a system. There are multiple approaches to assess the structural identifiability of a model including but not limited to a Taylor series expansion [7, 8, 9], the implicit function theorem [8, 9], generalizations of observability criteria [10, 11, 12], and differential algebra [^1] A subfield of algebra created by Ritt and his student Kolchin to study the existence of solutions to differential equations ([13, 14]). approaches [15, 16, 17].
@@
However, in practice, the quantity, frequency, and timing of available data have been shown to have large impacts on the reliability of parameter estimates. For examples in models of acute viral infections, see [2, 18, 19]. Furthermore, choices in the numerical discretizations (needed to approximate solutions) can also impact parameter estimates [20, 21]. Practical identifiability refers to whether parameters can be uniquely estimated given the available observations and the chosen parameter estimation method, and is closely related to the uncertainty quantification of model parameters. Some popular techniques to assess practical identifiability involve performing many replicate parameter estimations, which can become computationally expensive depending on the choice of estimation method and the complexity of the model.
@@
Weak-form based parameter estimation, which is typically much faster than output error-based methods, can address the challenge incurred by repeating the parameter estimation. The idea of using the weak form in system identification can be traced back to the Equations-of-Motion method described in [22] in which an equation (in the strong form) is multiplied by a compactly supported test function [^2] Shinbrot calls them method functions $\phi$ and then the equation is integrated. As the derivatives of the test functions are also assumed to have compact support, integration by parts removes the need to evaluate the derivatives of the data. Shinbrot proposed that the $\phi$’s be trigonometric functions with different frequencies, while Loeb and Cahen [23] later proposed piecewise polynomials (denoting their approach as the Modulating Functions Method). Recently, our group has developed a weak form-based parameter estimation and equation learning method in the spirit of Shinbrot, Loeb and Cahen, but with mathematically motivated test functions [24, 25, 26]. This choice of test functions improves both the computational efficiency of performing parameter estimation and is highly robust to noise in the data. These weak form based methods have also been applied to biological and environmental systems [27, 28, 29].
@@
One challenge in using this class of direct parameter estimation methods is that all variables must be measured. However, in biological systems, data is rarely observed for all model compartments. For example, epidemiological models are often calibrated to only a subset of model compartments, such as mortality or hospitalization data. Accordingly, researchers have considered differential elimination methods to collapse a system of low-order equations to an equivalent lower-dimensional high-order system with only observable variables [30, 31, 1, 32], and in some cases these systems have been cast into either integral or weak forms [1, 32]. However, previous applications of these elimination methods have not yielded a robust weak-form parameter estimation process.
@@
In this work, we propose a process for efficiently assessing a priori model identifiability by leveraging weak-form parameter estimation. First, we apply differential elimination methods to generate an input-output equation. At this step, we can also assess the structural identifiability of the model, which is prerequisite to practical identifiability. Second, we cast the input-output equation into its weak-form. Third, we implement a sampling-based method by performing weak-form parameter estimation on simulated data. This method is feasible even for large sample sizes and high levels of measurement error when the Weak form Estimation of Nonlinear Dynamics (WENDy) is used for parameter estimation, due to the method’s robustness to noise and computational efficiency. Finally, we assess practical identifiability from these samples using a newly defined criterion, $(e,q)-$identifiability.
@@
In Sections 2.1 and 2.2, we will give a brief overview of input-output-based methods to determine structural identifiability and methods used to determine practical identifiability. Then, in Section 2.3, we define $(e,q)-$identifiability, a practical identifiability criterion based on the noise in observed variables and the mean-square error for the parameter estimator. Next, in Section 3.1 we describe the generation of weak-form input-output and weak-form parameter estimation. These methods are applied together for two canonical biological examples in Section 3.2. Finally, in Section 4 we apply our definition of $(e,q)-$identifiability to the examples, compare this criterion to the average relative error, and lastly compare the computational efficiency of the weak-form parameter estimation to an output error method.
@@ 2 Background
In Sections 2.1 and 2.2, we provide background information on structural and practical identifiability. Readers already familiar with these topics are encouraged to skip to Section 2.3, in which, we introduce $(e,q)-$identifiability, a practical identifiability criterion based on relating the noise in observed variables and the mean-square error for the parameter estimator.
@@ 2.1 Structural Identifiability
Consider the system of ordinary differential equations $\dot{U}=\Theta(U,w)$, where $U$ is a vector of the state variables of the system and $w$ is a vector of model parameters. If the observations (output of the system) are given by $y(t)=\Omega(U,w)$, then structural identifiability can be defined as follows.
@@
One way of assessing the structural identifiability of a system of differential equations is by studying the input-output equation. Input-output equations have been used to study the structural identifiability of numerous biological systems, for examples, see [33, 34, 19, 35]. An input-output equation is a differential equation that depends only on the inputs, outputs, their respective derivatives, and the model parameters. Given an input-output equation for a system of differential equations, global structural identifiability can be restated as given below.
@@
A model is said to be globally structurally identifiable if coefficient map of the corresponding input-output equation, $c(w)$ is injective. [^3] Local structural identifiability can be similarly restated.
Input-output equations can be found from a system of differential equations using differential elimination algorithms, including the Ritt, the Ritt-Kolchin [14], and the Rosenfeld-Groebner [36] algorithms. There are many computational tools available to obtain input-output equations, including [17, 37, 38]. In this study, we use the DifferentialAlgebra v4 package [39] in Python to generate input-output equations, which applies a Rosenfeld-Groebner algorithm.
@@ 2.2 Practical Identifiability
There are many proposed methods to assess the practical identifiability of a system, and there is no commonly agreed-upon single criterion to determine practical identifiability. The Fisher Information Matrix (FIM) can be used to approximate the parameter covariance matrix and as a measure of local practical identifiability [40, 41], however, this criterion is not always generalizability to nonlinear systems. The FIM method has been used to assess the practical identifiability in biological systems [33, 3]. Profile Likelihood based methods test practical identifiability through successive refittings in the parameter space, which give insight into the shape of the likelihood and whether it is uniquely maximized [42]. Profile likelihood methods have been used to assess practical identifiability in biological systems as well [43, 35]. However, methods, like Profile Likelihood, rely on the special relationship between the output error objective function and the maximum likelihood.
Sampling-based methods of assessing practical identifiability involve repeated estimation performed using a prescribed error model for the observed data. From these repeated estimations, one criterion that has been proposed to assess practical identifiability is the average relative error. Examples of this criterion applied to biological systems include [44, 19]. Sampling-based methods can be applied to any choice of objective function, but they also require numerous repetitions of the chosen parameter estimation method, which can become challenging to implement if the estimation method is not computationally efficient.
@@
We are interested in using a sampling-based method of assessing practical identifiability. By choosing such a method, we can avoid making assumptions about the local linearization of the system, which is required to find the Fisher Information Matrix. And, we do not need to restrict our choice of objective function, which is needed for Profile Likelihood Analysis. While sampling-based methods can pose a computational challenge, recent advances in weak-form parameter estimation can be leveraged to perform the many required repetitions of the estimation quickly.
We propose a practical identifiability criterion that depends on the mean-squared error and a fixed level of additive error in the data. Below we define the $\underline{(\mathbf{e,q})-}$ identifiablity, which can be computed for systems and data using either output error or equation error optimization.
@@
$y(t)=\Omega(U,w)+\epsilon,$
where $\epsilon\sim\mathcal{F}$ and $\mathcal{F}$ is an arbitrary distribution with variance $\sigma^{2}$.
@@
Let $e$ denote the additive error ratio such that $\sigma=e\text{RMS}(\Omega(t))$, where RMS is the root mean square, i.e., $\text{RMS}(\Omega(t))=\sqrt{\frac{1}{T}\int_{0}^{T}\Omega(t)^{2}}$.
@@
Let $q$ denote the estimate error ratio such that $M_{i}=(q|w_{i}|)^{2}$, where $M_{i}$ is the maximum tolerated value of some error measurement in the parameter estimator, $\widehat{w}_{i}$.
Note that in the case that the true parameter value is unknown, we let $M_{i}=\mathbb{E}(\widehat{w}_{i_{k}})$. Finally, we define an identifiability criterion relating these two ratios.
@@
A parameter is said to be $\underline{(\mathbf{e,q})}$-identifiable if, for a given additive error ratio, $e$, the mean-squared error (MSE) of the parameter estimator, $\widehat{w}_{i}$, is below $M_{i}$, i.e., $MSE(\widehat{w}_{i}(\mathbf{e}))<(\mathbf{q}|w_{i}|)^{2}~{}.$ Furthermore, a model is said to be $\underline{(\mathbf{e,q})}$-identifiable if the above holds for all $i$.
@@
Here, $\text{MSE}(\widehat{w})=\mathbb{E}[(\widehat{w}_{k}-w)^{2}]=\text{Var}(\widehat{w})+\text{Bias}(\widehat{w},w)^{2}$. Since the MSE can be decomposed into the variance and squared bias, $(e,q)-$indentifiability indicates that we can obtain parameter estimates that are both near the true parameter and near each other with respect to a cut-off, $(q|w|)^{2}$, that is scaled by the size of the parameter and the choice of acceptable estimate error ratio $q$.
3 Weak-form Input-output equations
In this section, we combine established differential elimination methods with recent advances in weak-form parameter estimation to accurately estimate parameters from systems with unobserved compartments. In Section 3.1, we describe a method to generate weak-form input-output equations by using differential elimination, then briefly present a method of parameter estimation using these equations. Lastly, in Section 3.2, we illustrate the application of these methods to two canonical biological systems.
@@ 3 Weak-form Input-output equations
3.1 Weak-form parameter estimation for systems with unobserved variables

### v3

@@ Abstract
In this work, we define a practical identifiability criterion, $(e,q)$-identifiability, based on a parameter $e$, reflecting the noise in observed variables, and a parameter $q$, reflecting the mean-square error of the parameter estimator. This criterion is better able to encompass changes in the quality of the parameter estimate(s) due to increased noise in the data (compared to existing criteria based solely on average relative errors). We illustrate the usefulness of the criteria in several challenging identifiability studies, involving parameter estimation in partially observed systems. Furthermore, we leverage a weak-form equation error-based method of parameter estimation for systems with unobserved variables to assess practical identifiability far more quickly in comparison to output error-based parameter estimation. We do so by generating weak-form input-output equations using differential algebra techniques, as previously proposed by Boulier et al [1], and then applying Weak form Estimation of Nonlinear Dynamics (WENDy) to obtain parameter estimates. This method is computationally efficient and robust to noise, as demonstrated through two classical biological modeling examples.
@@ 1 Introduction
Parameter estimation is central to mathematical modeling in the biological sciences. However, despite the increasing sophistication of computational and statistical techniques over the years, there are many examples of researchers using the same model and obtaining widely varying parameter estimates for the same biological phenomena [2, 3, 4, 5, 6]. It is well-recognized that parameter estimation in dynamical systems remains a challenging task [7]. While some of this variation is due to causes such as quality and quantity of data and uncertainty in initial conditions, the primary reason for that difficulty lies in the relationship between the model structure and data. Analyzing model identifiability, i.e., the ability to recover unique parameters of a system from observations, is thus critical to developing robust and replicable inference in biological systems.
@@
Structural identifiability addresses whether model parameters can be uniquely determined given the observation of some output variables of a system. There are multiple approaches to assess the structural identifiability of a model including but not limited to a Taylor series expansion [8, 9, 10], the implicit function theorem [9, 10], generalizations of observability criteria [11, 12, 13], and differential algebra [^1] A subfield of algebra created by Ritt and his student Kolchin to study the existence of solutions to differential equations ([14, 15]). approaches [16, 17, 18, 19].
@@
However, in practice, the quantity, frequency, and timing of available data have been shown to have large impacts on the reliability of parameter estimates. For examples in models of acute viral infections, we direct the reader to [2, 20, 21]. Furthermore, choices in the numerical discretizations (needed to approximate solutions) can also impact parameter estimates [22, 23]. Practical identifiability [^2] This has also been referred to as statistical identifiability. refers to whether parameters can be uniquely estimated given the available observations and the chosen parameter estimation method, including numerics, and is directly related to uncertainty quantification of model parameters. Practical identifiability methods can be applied to study estimator properties for a model with a specific observed dataset (a posteriori), or to study general properties of a model in different hypothetical data scenarios, i.e., noise and collection frequency (a priori). The latter can provide important insights to guide experimental design, power analyses, and data collection.
@@
While simple models can benefit from traditional statistical methods, such as the Fisher Information Matrix, most practical identifiability techniques for complex models rely on repeated simulations, which can become computationally expensive. In contrast, weak-form based parameter estimation, which is typically much faster than output error-based methods, can address the challenge incurred by repeating the parameter estimation. The idea of using the weak form in system identification can be traced back to the Equations-of-Motion method described in [24], in which an equation (in the strong form) is multiplied by a compactly supported test function [^3] Shinbrot calls them method functions $\phi$ and then the equation is integrated. As the derivatives of the test functions are also assumed to have compact support, integration by parts removes the need to evaluate the derivatives of the data. Shinbrot proposed that the $\phi$’s be trigonometric functions with different frequencies, while Loeb and Cahen [25] later proposed piecewise polynomials (denoting their approach as the Modulating Functions Method). Recently, our group has developed a suite of weak form-based parameter estimation and equation learning methods (WENDy and WSINDy) in the spirit of Shinbrot, Loeb and Cahen, but with mathematically motivated test functions [26, 27, 28]. This choice of test functions improves both the computational efficiency of performing parameter estimation and is highly robust to noise in the data. These weak form-based methods have also been applied to biological and environmental systems [29, 30, 31].
@@
One challenge in using WENDy type weak-form estimation methods is that, thus far, all state variables had to be measured. However, in biological systems, data is rarely observed for all model compartments. For example, epidemiological models are often calibrated to only a subset of model compartments, such as mortality or hospitalization data. Accordingly, in this study, we consider differential elimination methods to collapse a system of low-order equations to an equivalent lower-dimensional high-order system with only observable variables. This technique has been applied in [32, 33, 1, 34], and, in some cases, these systems have been cast into either integral or weak forms [1, 34]. However, previous applications of these elimination methods have not yielded a robust weak-form parameter estimation process.
@@
In this work, we propose a process for efficiently assessing a priori model identifiability by leveraging weak-form parameter estimation. First, we apply differential elimination methods to generate an input-output equation, for the states which will be observed. At this step, we can also assess the structural identifiability of the model, which is a prerequisite to practical identifiability. Second, we cast the input-output equation into its weak form. Third, we perform weak-form parameter estimation on repeated simulations of noisy data. Note, unlike output-error based methods, this method is feasible even for large sample sizes and high levels of measurement error when the Weak form Estimation of Nonlinear Dynamics (WENDy) is used for parameter estimation, due to the method’s robustness to noise and computational efficiency. Finally, we assess practical identifiability from these samples using a newly defined criterion, which we refer to as $(e,q)$-identifiability.
@@
In Sections 2.1 and 2.2, we will give a brief overview of input-output-based methods to determine structural identifiability and methods used to determine practical identifiability. Then, in Section 2.3, we define $(e,q)$-identifiability, a practical identifiability criterion based on the noise in observed variables and the mean-square error for the corresponding parameter estimator. Next, in Sections 3.1 and 3.2 we describe the generation of weak-form input-output equations and weak-form parameter estimation. The structural identifiability and limitations of the weak-form input-output parameter estimation are discussed in 3.3. The weak-form input-ouput parameter estimation is applied to two canonical biological examples in Section 3.4. Finally, in Section 4 we apply our definition of $(e,q)$-identifiability to the examples, compare this criterion to the average relative error, and lastly compare the computational efficiency of the weak-form parameter estimation to an output error method.
@@ 2 Background
In Sections 2.1 and 2.2, we provide background information on structural and practical identifiability. Readers already familiar with these topics are encouraged to skip to Section 2.3, in which we introduce $(e,q)$-identifiability, a practical identifiability criterion based on relating the noise in observed variables and the mean-square error for the parameter estimator.
@@ 2.1 Structural Identifiability
Consider the system of ordinary differential equations $\dot{U}=\Theta(U,w)$, where $U$ is a vector of the state variables of the system and $w$ is a vector of model parameters. If the noiseless observations (output of the system) are given by $y(t)=\Omega(U,w)$, then structural identifiability can be defined as follows.
@@
Otherwise, a model is said to be structurally unidentifiable.
One way of assessing the structural identifiability of a system of differential equations is by studying the input-output equation. Input-output equations have been used to study the structural identifiability of numerous biological systems, for examples, see [35, 36, 21, 37, 38]. An input-output equation is a differential equation that depends only on the inputs, outputs, their respective derivatives, and the model parameters. Given an input-output equation for a system of differential equations, global structural identifiability can be restated as given below.
@@
A model is said to be globally structurally identifiable if coefficient map of the corresponding input-output equation, $c(w)$ is injective. [^4] Local structural identifiability can be similarly restated. [^5] For this definition to be equivalent to Definition 1, there is an criterion on the Wronskian of the differential monomials from the input-output equation. For more details, see [39].
Input-output equations can be found from a system of differential equations using differential elimination algorithms, including the Ritt, the Ritt-Kolchin [15], and the Rosenfeld-Groebner [40] algorithms, or with a more recently developed, projection-based algorithm [19]. There are many computational tools available to obtain input-output equations, including [18, 41, 42]. In this study, we use the DifferentialAlgebra v4 package [43] in Python to generate input-output equations, which applies a Rosenfeld-Groebner algorithm. Note, an input-output equation may not exist or may not be computationally feasible to discover; however, addressing the weak-form estimation in such cases is a subject of future work. This paper focuses on a weak-form method for the case that an input-output equation, as referred to in Definition 5, can be found.
@@ 2.2 Practical Identifiability
There are many proposed methods to assess the practical identifiability of a system, and there is no commonly agreed-upon single criterion to determine practical identifiability. The Fisher Information Matrix (FIM) can be used to approximate the parameter covariance matrix and as a measure of local practical identifiability [44, 45], however, this criterion is difficult to calculate in more complex systems where likelihood modes may be difficult to find. The FIM method has been used to assess the practical identifiability in biological systems [35, 3]. In these more complicated systems, Profile Likelihood based methods test practical identifiability through successive refittings in the parameter space, which give insight into the shape of the likelihood and whether it is uniquely maximized [^6] Depending on the choice of estimator or proxy used for the likelihood, this problem may be restated as a minimization. [46]. Profile likelihood methods have been used to assess practical identifiability in biological systems as well [47, 37].
Simulation-based methods of assessing practical identifiability involve repeated estimation performed using a prescribed error model for the observed data. From these repeated estimations, one criterion that has been proposed to assess practical identifiability is the average relative error. Examples of applications of this criterion applied to biological systems include [48, 21]. Simulation-based methods can be applied to any choice of objective function, but they also require numerous repetitions of the chosen parameter estimation method, which can become challenging to implement if the estimation method is not computationally efficient.
@@
We are interested in using a simulation-based method of assessing practical identifiability. By choosing such a method, we can avoid making assumptions about the local approximation of the system, which is required to find the Fisher Information Matrix. Likewise, we do not need to restrict our choice of objective function, as in Profile Likelihood Analysis. While simulation-based methods can pose a computational challenge, recent advances in weak-form parameter estimation can be leveraged to perform the many required repetitions of the estimation quickly.
We propose a practical identifiability criterion that depends on the mean-squared error on the chosen estimator. Below we define the ($e,q$)-identifiability, which can be computed for systems and data using either output error or equation error optimization; however, in this paper we use equation error optimization.
@@
$y(t)={\color[rgb]{0,0,0}\mathcal{Y}(\Omega(U{\color[rgb]{0,0,0}(t)},w),\epsilon)}$
where $\epsilon\sim\mathcal{F}$ and $\mathcal{F}$ is an arbitrary distribution with variance $\sigma^{2}$. For example, in the case of additive noise, $y(t)=\Omega(U{\color[rgb]{0,0,0}(t)},w)+\epsilon$, and, in the case of multiplicative noise, $y(t)=\epsilon\Omega(U{\color[rgb]{0,0,0}(t)},w)$.
Note that $\Omega$ is a vector of the observation of states in the absence of noise as described in Section 2.1. The examples in this study and the following definitions are written in terms of a single observation state. While $(e,q)$- identifiability can be applied for more than one observation state, we leave choosing an appropriate scaling of multiple observation states for future work.
@@
Let $e$ denote the observation error ratio such that $e=\sigma/\text{RMS}(\Omega)$, where $\sigma^{2}$ is variance of the error distribution, as defined above, and RMS is the root mean square, i.e., $\text{RMS}(\Omega)={\frac{1}{\sqrt{T_{m}-T_{0}}}\int_{T_{0}}^{T_{M}}\Omega(t)^{2}dt}$. Here, we assume that observations are made at time points $\{T_{0},...,T_{M}\}$.
@@
Let $q$ denote the estimator error ratio such that $q=\sqrt{M_{i}}/|w_{i}|$, where $M_{i}$ is the maximum tolerated mean-squared error in the parameter estimator, $\widehat{w}_{i}$. Here, $\text{MSE}(\widehat{w})=\mathbb{E}[(\widehat{w}_{k}-w)^{2}]=\text{Var}(\widehat{w})+\text{Bias}(\widehat{w},w)^{2}$, and $w$ is the true value of the parameter.
Finally, we define an identifiability criterion relating these two ratios.
@@
A parameter, $w_{i}$, is said to be $(e,q)$-identifiable if, for a given observation error ratio, $e$, the mean-squared error (MSE) of the parameter estimator, $\widehat{w}_{i}$, is below the maximum tolerated mean-squared error, $M_{i}$, i.e., $MSE(\widehat{w}_{i}(\mathbf{e}))<(\mathbf{q}{\color[rgb]{0,0,0}w_{i}})^{2}~.$ Furthermore, a model is said to be $(e,q)$-identifiable if the above holds for all $i$.
@@
Since the MSE can be decomposed into the variance and squared bias, $(e,q)$-identifiability indicates that we can obtain parameter estimates that are both near the true parameter and near each other, in other words both accurate and precise, with respect to a cut-off, $(qw)^{2}$, that is scaled by the size of the parameter and the choice of acceptable estimate error ratio $q$. Note, $(e,q)$-identifiability could be applied a priori or a posteriori. When the true parameter value is unknown, as will be the case in most a posteriori applications, and when we have an unbiased estimator, we can let $w_{i}=\mathbb{E}(\widehat{w}_{i_{k}})$.
As with most simulation-based tests of practical identifiability, our goal in using the values $e$ and $q$ is to relate the uncertainty in the parameter estimator to uncertainty in the observations, and for these values to be interpretable and flexible in application. The observation error ratio, $e$, represents a ratio of a measurement ($RMS$) of the observation magnitude and the variance of the error, which allows us to use the criterion independently of the scale of the data. The estimator error ratio, $q$, represents the ratio of parameter magnitude and mean-squared error of the parameter estimator, which allows the use of the criterion for a variety of ODE systems and estimators. For example, if two models are (5,20)-identifiable, i.e., $e=5\%$ and $q=20\%$, for some set of data, then for variance in the error of the data that is $5\%$ of the average observation, we expect all parameter MSEs to remain below $20\%$ of the magnitude of the true parameters. Thus, while these models may have different scalings, they can still be compared directly.
@@
3 Weak-form parameter estimation for systems with unobserved variables
In this section, we combine established differential elimination methods with recent advances in weak-form parameter estimation to accurately estimate parameters from systems with unobserved compartments. In Section 3.1, we describe a method to generate weak-form input-output equations by using differential elimination, then in Section 3.2 present a method of parameter estimation using these equations. Next, in Section 3.3, we discuss structural identifiability and limitations of the weak-form input-output parameter estimation. Lastly, in Section 3.4, we illustrate the application of these methods to two canonical biological systems.
3.1 Weak-form Input-output equations